# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ Appendix G Comparison of schedules
Table 11: Penalised return statistics comparing the schedule with restart schedules to the traditional linear schedule under the short runs (200-500 episodes $\times$ maximal number of time steps) and the long runs (2000-5000 episodes $\times$ maximal number of time steps). Bold indicates the top performance and any additional algorithms that are within one pooled standard error.
Domain Algorithm $R_{\text{pen}}^{\pm}$ (signed) $R_{\text{pen}}$ (positive)
Mean $\pm$ SE Min Mean $\pm$ SE Min
Cartpole (short runs) MDPO-Robust-Lag (linear) 107.8 $\pm$ 4.4 64.0 33.8 $\pm$ 1.2 24.4
MDPO-Robust-Lag (restart) 106.9 $\pm$ 4.7 57.9 34.3 $\pm$ 1.1 25.3
MDPO-Robust-Lag (geometric) 108.2 $\pm$ 4.5 64.3 34.2 $\pm$ 1.2 24.4
MDPO-Robust-Augmented-Lag (linear) 108.1 $\pm$ 4.7 60.5 34.4 $\pm$ 1.2 25.0
MDPO-Robust-Augmented-Lag (restart) 107.6 $\pm$ 4.3 68.1 33.8 $\pm$ 1.2 24.0
MDPO-Robust-Augmented-Lag (geometric) 102.3 $\pm$ 3.9 64.6 32.3 $\pm$ 0.9 24.2
Cartpole (long runs) MDPO-Robust-Lag (linear) 111.4 $\pm$ 7.5 36.8 38.6 $\pm$ 1.7 16.1
MDPO-Robust-Lag (restart) 109.0 $\pm$ 6.7 39.0 37.3 $\pm$ 1.9 12.3
MDPO-Robust-Lag (geometric) 107.6 $\pm$ 7.2 32.6 37.7 $\pm$ 1.8 13.1
MDPO-Robust-Augmented-Lag (linear) 108.5 $\pm$ 5.9 48.3 36.3 $\pm$ 1.6 19.1
MDPO-Robust-Augmented-Lag (restart) 106.8 $\pm$ 6.4 42.0 36.5 $\pm$ 1.8 13.1
MDPO-Robust-Augmented-Lag (geometric) 106.2 $\pm$ 6.6 36.0 36.4 $\pm$ 1.9 13.3
Inventory Management (short runs) MDPO-Robust-Lag (linear) -1466.7 $\pm$ 30.0 -1616.6 -1467.1 $\pm$ 29.5 -1616.6
MDPO-Robust-Lag (restart) -1081.1 $\pm$ 52.0 -1337.8 -1082.5 $\pm$ 50.7 -1337.8
MDPO-Robust-Lag (geometric) -1194.1 $\pm$ 32.1 -1352.6 -1194.7 $\pm$ 31.3 -1352.6
MDPO-Robust-Augmented-Lag (linear) -1383.5 $\pm$ 45.6 -1585.6 -1384.1 $\pm$ 44.8 -1585.6
MDPO-Robust-Augmented-Lag (restart) -837.4 $\pm$ 21.2 -939.7 -838.1 $\pm$ 20.3 -939.7
MDPO-Robust-Augmented-Lag (geometric) -1258.4 $\pm$ 40.5 -1434.6 -1260.4 $\pm$ 38.0 -1434.6
Inventory Management (long runs) MDPO-Robust-Lag (linear) -837.3 $\pm$ 21.3 -939.1 -838.0 $\pm$ 20.3 -939.1
MDPO-Robust-Lag (restart) -931.2 $\pm$ 32.0 -1117.7 -932.5 $\pm$ 30.6 -1117.7
MDPO-Robust-Lag (geometric) -837.3 $\pm$ 21.4 -939.2 -838.0 $\pm$ 20.4 -939.2
MDPO-Robust-Augmented-Lag (linear) -837.3 $\pm$ 21.4 -938.9 -838.0 $\pm$ 20.3 -938.9
MDPO-Robust-Augmented-Lag (restart) -837.4 $\pm$ 21.3 -939.7 -838.0 $\pm$ 20.5 -939.7
MDPO-Robust-Augmented-Lag (geometric) -837.3 $\pm$ 21.5 -939.2 -838.1 $\pm$ 20.3 -939.2
3-D Inventory Management (short runs) MDPO-Robust-Lag (linear) -832.6 $\pm$ 441.0 -7278.5 -5011.3 $\pm$ 213.8 -8576.3
MDPO-Robust-Lag (restart) -1112.1 $\pm$ 444.2 -7558.3 -5353.0 $\pm$ 216.9 -8926.7
MDPO-Robust-Lag (geometric) -3096.4 $\pm$ 449.0 -9662.6 -6051.9 $\pm$ 254.2 -10256.4
MDPO-Robust-Augmented-Lag (linear) -688.5 $\pm$ 445.4 -7202.3 -4951.8 $\pm$ 205.8 -8471.1
MDPO-Robust-Augmented-Lag (restart) -1725.5 $\pm$ 443.2 -8179.6 -5450.3 $\pm$ 220.0 -9136.5
MDPO-Robust-Augmented-Lag (geometric) -2863.4 $\pm$ 449.0 -9571.2 -5822.4 $\pm$ 256.6 -10229.3
3-D Inventory Management (long runs) MDPO-Robust-Lag (linear) 1596.9 $\pm$ 387.2 -3950.6 -1727.3 $\pm$ 176.4 -4888.8
MDPO-Robust-Lag (restart) 1100.4 $\pm$ 358.3 -3951.0 -2198.4 $\pm$ 198.4 -5525.9
MDPO-Robust-Lag (geometric) 1652.3 $\pm$ 374.5 -3607.3 -1518.4 $\pm$ 176.0 -4544.3
MDPO-Robust-Augmented-Lag (linear) 5669.4 $\pm$ 353.2 939.5 -1294.3 $\pm$ 137.0 -3591.4
MDPO-Robust-Augmented-Lag (restart) 5683.9 $\pm$ 346.5 902.2 -928.3 $\pm$ 128.2 -3117.7
MDPO-Robust-Augmented-Lag (geometric) 7172.7 $\pm$ 332.9 2773.7 -1212.7 $\pm$ 145.7 -3895.8

### v2

@@ Appendix H Comparison of schedules
Table 11: Return and penalised return statistics comparing per-epoch schedules for the penalty parameter $\alpha$ under the short runs (200-500 episodes $\times$ maximal number of time steps) and the long runs (2000-5000 episodes $\times$ maximal number of time steps).
Algorithm Return $R_{\text{pen}}^{\pm}$ (signed) $R_{\text{pen}}$ (positive)
Mean $\pm$ SE Min Mean $\pm$ SE Min Mean $\pm$ SE Min
Cartpole (short runs)
MDPO-Robust (fixed) 63.1 $\pm$ 0.1 56.4 -229.7 $\pm$ 179.6 -2610.6 -294.1 $\pm$ 171.4 -2610.8
MDPO-Robust (linear) 62.9 $\pm$ 0.1 55.0 -482.5 $\pm$ 182.3 -2487.1 -514.6 $\pm$ 177.4 -2487.1
MDPO-Robust (restart) 63.1 $\pm$ 0.1 55.2 -252.8 $\pm$ 179.1 -2520.5 -304.8 $\pm$ 172.0 -2520.6
MDPO-Robust (geometric) 63.1 $\pm$ 0.1 55.8 -148.0 $\pm$ 173.5 -2480.3 -219.3 $\pm$ 164.7 -2481.5
MDPO-Robust-Lag (fixed) 34.4 $\pm$ 0.3 24.9 108.4 $\pm$ 4.4 65.9 34.2 $\pm$ 1.2 24.9
MDPO-Robust-Lag (linear) 35.3 $\pm$ 0.3 25.2 107.1 $\pm$ 5.0 53.0 34.5 $\pm$ 1.2 25.0
MDPO-Robust-Lag (restart) 34.9 $\pm$ 0.3 25.7 107.2 $\pm$ 4.9 53.1 34.3 $\pm$ 1.1 25.4
MDPO-Robust-Lag (geometric) 34.8 $\pm$ 0.3 25.0 109.4 $\pm$ 4.8 60.2 34.5 $\pm$ 1.2 25.0
MDPO-Robust-Augmented-Lag (fixed) 34.6 $\pm$ 0.3 24.7 110.8 $\pm$ 4.9 61.6 34.6 $\pm$ 1.2 24.7
MDPO-Robust-Augmented-Lag (linear) 35.2 $\pm$ 0.3 25.4 106.6 $\pm$ 5.0 52.6 34.2 $\pm$ 1.2 22.9
MDPO-Robust-Augmented-Lag (restart) 35.4 $\pm$ 0.3 25.7 107.1 $\pm$ 5.4 45.6 34.4 $\pm$ 1.3 21.7
MDPO-Robust-Augmented-Lag (geometric) 35.1 $\pm$ 0.3 24.6 109.7 $\pm$ 5.1 56.4 34.7 $\pm$ 1.2 24.6
Cartpole (long runs)
MDPO-Robust (fixed) 63.2 $\pm$ 0.1 56.1 -109.7 $\pm$ 176.0 -2647.1 -185.5 $\pm$ 167.1 -2648.3
MDPO-Robust (linear) 33.7 $\pm$ 0.1 29.2 -145.3 $\pm$ 24.1 -526.4 -165.0 $\pm$ 23.0 -541.6
MDPO-Robust (restart) 63.1 $\pm$ 0.1 55.8 -173.6 $\pm$ 174.9 -2537.4 -237.1 $\pm$ 167.0 -2538.1
MDPO-Robust (geometric) 63.2 $\pm$ 0.1 56.1 -120.4 $\pm$ 179.5 -2662.1 -193.7 $\pm$ 170.7 -2662.7
MDPO-Robust-Lag (fixed) 41.6 $\pm$ 0.2 32.1 105.1 $\pm$ 9.5 9.6 38.6 $\pm$ 2.5 2.6
MDPO-Robust-Lag (linear) 32.3 $\pm$ 0.3 22.9 72.0 $\pm$ 4.1 38.6 15.4 $\pm$ 1.8 -4.8
MDPO-Robust-Lag (restart) 41.8 $\pm$ 0.2 32.2 107.6 $\pm$ 9.3 14.5 39.2 $\pm$ 2.3 5.8
MDPO-Robust-Lag (geometric) 41.4 $\pm$ 0.2 31.7 103.7 $\pm$ 9.4 8.0 38.6 $\pm$ 2.5 1.3
MDPO-Robust-Augmented-Lag (fixed) 41.2 $\pm$ 0.2 31.4 106.2 $\pm$ 9.1 13.7 38.6 $\pm$ 2.2 5.0
MDPO-Robust-Augmented-Lag (linear) 30.0 $\pm$ 0.3 21.5 32.8 $\pm$ 2.8 12.2 -18.8 $\pm$ 1.3 -34.6
MDPO-Robust-Augmented-Lag (restart) 40.5 $\pm$ 0.2 30.5 108.7 $\pm$ 8.2 24.1 38.3 $\pm$ 1.9 10.7
MDPO-Robust-Augmented-Lag (geometric) 40.9 $\pm$ 0.2 31.4 102.1 $\pm$ 8.9 9.2 37.6 $\pm$ 2.7 -2.1
Inventory Management (short runs)
MDPO-Robust (fixed) 118.0 $\pm$ 1.8 72.3 -3306.4 $\pm$ 9.7 -3350.0 -3306.4 $\pm$ 9.7 -3350.0
MPDO-Robust (linear) 98.4 $\pm$ 5.7 -59.6 -2764.0 $\pm$ 20.5 -2808.5 -2764.2 $\pm$ 20.2 -2808.5
MPDO-Robust (restart) 105.1 $\pm$ 6.3 -20.2 -3404.8 $\pm$ 32.9 -3634.3 -3404.8 $\pm$ 32.9 -3634.3
MPDO-Robust (geometric) 114.6 $\pm$ 2.3 49.7 -3166.4 $\pm$ 6.9 -3207.6 -3166.5 $\pm$ 6.9 -3207.6
MPDO-Robust-Lag (fixed) 60.7 $\pm$ 10.7 -228.5 -930.9 $\pm$ 32.3 -1118.2 -932.2 $\pm$ 30.8 -1118.2
MPDO-Robust-Lag (linear) 93.2 $\pm$ 4.7 -55.0 -2186.8 $\pm$ 6.9 -2248.8 -2187.3 $\pm$ 6.4 -2248.8
MPDO-Robust-Lag (restart) 88.3 $\pm$ 4.8 -61.3 -1880.7 $\pm$ 16.5 -2003.7 -1881.1 $\pm$ 16.1 -2003.7
MPDO-Robust-Lag (geometric) 67.2 $\pm$ 9.0 -156.8 -1046.3 $\pm$ 10.2 -1122.4 -1048.1 $\pm$ 8.0 -1122.4
MPDO-Robust-Augmented-Lag (fixed) 60.9 $\pm$ 10.6 -221.4 -938.6 $\pm$ 34.4 -1172.2 -940.1 $\pm$ 32.9 -1172.2
MPDO-Robust-Augmented-Lag (linear) 91.8 $\pm$ 4.4 -47.8 -2033.0 $\pm$ 8.0 -2054.6 -2033.3 $\pm$ 7.6 -2054.6
MPDO-Robust-Augmented-Lag (restart) 85.6 $\pm$ 5.9 -94.9 -1902.5 $\pm$ 6.5 -1954.2 -1902.7 $\pm$ 6.2 -1954.2
MPDO-Robust-Augmented-Lag (geometric) 60.0 $\pm$ 9.3 -206.9 -1095.9 $\pm$ 18.3 -1128.7 -1097.2 $\pm$ 16.3 -1128.7
Inventory Management (long runs)
MDPO-Robust (fixed) 119.7 $\pm$ 1.8 72.4 -3416.7 $\pm$ 4.1 -3429.7 -3416.7 $\pm$ 4.1 -3429.7
MPDO-Robust (linear) 114.8 $\pm$ 2.3 53.6 -3166.7 $\pm$ 6.8 -3208.2 -3166.7 $\pm$ 6.8 -3208.2
MPDO-Robust (restart) 119.9 $\pm$ 1.8 74.4 -3425.2 $\pm$ 5.4 -3460.2 -3425.2 $\pm$ 5.4 -3460.2
MPDO-Robust (geometric) 114.7 $\pm$ 2.3 51.0 -3166.5 $\pm$ 6.9 -3207.6 -3166.6 $\pm$ 6.9 -3207.6
MPDO-Robust-Lag (fixed) 66.9 $\pm$ 7.1 -166.5 -837.4 $\pm$ 21.4 -939.3 -838.0 $\pm$ 20.4 -939.3
MPDO-Robust-Lag (linear) 66.8 $\pm$ 7.1 -165.7 -837.3 $\pm$ 21.4 -939.1 -838.0 $\pm$ 20.3 -939.1
MPDO-Robust-Lag (restart) 66.8 $\pm$ 7.2 -169.1 -837.3 $\pm$ 21.6 -939.8 -838.0 $\pm$ 20.5 -939.8
MPDO-Robust-Lag (geometric) 66.8 $\pm$ 7.2 -168.9 -837.3 $\pm$ 21.6 -939.7 -837.9 $\pm$ 20.6 -939.7
MPDO-Robust-Augmented-Lag (fixed) 66.8 $\pm$ 7.1 -167.7 -837.2 $\pm$ 21.5 -939.1 -837.9 $\pm$ 20.5 -939.1
MPDO-Robust-Augmented-Lag (linear) 66.8 $\pm$ 7.1 -163.4 -837.4 $\pm$ 21.3 -939.2 -838.0 $\pm$ 20.3 -939.2
MPDO-Robust-Augmented-Lag (restart) 66.8 $\pm$ 7.1 -162.8 -837.4 $\pm$ 21.2 -939.0 -837.9 $\pm$ 20.4 -939.0
MPDO-Robust-Augmented-Lag (geometric) 66.9 $\pm$ 7.1 -162.6 -837.4 $\pm$ 21.2 -939.2 -838.1 $\pm$ 20.2 -939.2
3-D Inventory Management (short runs)
MDPO-Robust (fixed) 54.9 $\pm$ 0.4 -9.6 -7956.9 $\pm$ 457.5 -14747.8 -9376.8 $\pm$ 330.2 -14836.3
MPDO-Robust (linear) 56.5 $\pm$ 0.4 -8.1 -8716.0 $\pm$ 458.1 -15404.0 -9918.5 $\pm$ 343.1 -15469.0
MPDO-Robust (restart) 56.3 $\pm$ 0.4 -8.7 -8693.9 $\pm$ 455.4 -15503.7 -9866.2 $\pm$ 333.9 -15531.6
MPDO-Robust (geometric) 53.6 $\pm$ 0.4 -11.5 -7337.6 $\pm$ 459.1 -14074.2 -8846.6 $\pm$ 315.8 -14173.9
MPDO-Robust-Lag (fixed) 41.4 $\pm$ 0.5 -24.6 -1416.9 $\pm$ 446.1 -7726.9 -5088.9 $\pm$ 221.7 -8560.6
MPDO-Robust-Lag (linear) 41.8 $\pm$ 0.5 -25.5 -1584.6 $\pm$ 447.5 -8169.0 -5178.0 $\pm$ 222.3 -8982.4
MPDO-Robust-Lag (restart) 41.9 $\pm$ 0.5 -24.9 -1648.0 $\pm$ 446.1 -8138.0 -5308.1 $\pm$ 223.1 -8995.2
MPDO-Robust-Lag (geometric) 41.1 $\pm$ 0.5 -28.2 -1278.3 $\pm$ 449.6 -7571.7 -5060.9 $\pm$ 219.1 -8659.8
MPDO-Robust-Augmented-Lag (fixed) 39.5 $\pm$ 0.5 -29.6 -525.0 $\pm$ 441.2 -6861.9 -4957.0 $\pm$ 211.2 -8289.3
MPDO-Robust-Augmented-Lag (linear) 42.4 $\pm$ 0.5 -23.7 -1929.8 $\pm$ 446.9 -8420.5 -5311.8 $\pm$ 228.7 -9203.4
MPDO-Robust-Augmented-Lag (restart) 40.7 $\pm$ 0.5 -26.8 -1070.1 $\pm$ 441.9 -7347.8 -5155.6 $\pm$ 216.3 -8671.2
MPDO-Robust-Augmented-Lag (geometric) 42.8 $\pm$ 0.5 -25.1 -2123.5 $\pm$ 450.2 -8567.3 -5507.8 $\pm$ 228.9 -9303.0
3-D Inventory Management (long runs)
MDPO-Robust (fixed) 87.0 $\pm$ 0.4 24.4 -23312.6 $\pm$ 525.6 -31314.1 -23459.7 $\pm$ 499.3 -31314.1
MPDO-Robust (linear) 96.3 $\pm$ 0.4 36.0 -27752.5 $\pm$ 547.7 -35936.5 -27902.6 $\pm$ 525.3 -35936.5
MPDO-Robust (restart) 90.8 $\pm$ 0.4 27.4 -25169.7 $\pm$ 523.8 -32979.5 -25265.0 $\pm$ 504.2 -32979.5
MPDO-Robust (geometric) 82.2 $\pm$ 0.4 20.0 -21022.5 $\pm$ 487.2 -28373.0 -21098.2 $\pm$ 469.0 -28373.0
MPDO-Robust-Lag (fixed) 32.8 $\pm$ 0.5 -32.4 2617.9 $\pm$ 378.6 -2601.7 -1011.4 $\pm$ 169.9 -3767.4
MPDO-Robust-Lag (linear) 33.2 $\pm$ 0.5 -31.9 2538.8 $\pm$ 368.0 -2793.7 -1353.1 $\pm$ 197.6 -4804.2
MPDO-Robust-Lag (restart) 34.0 $\pm$ 0.4 -31.2 2062.3 $\pm$ 375.3 -3333.9 -1317.1 $\pm$ 191.5 -4687.1
MPDO-Robust-Lag (geometric) 31.7 $\pm$ 0.5 -34.9 3269.5 $\pm$ 371.3 -2036.2 -1097.3 $\pm$ 173.3 -4299.3
MPDO-Robust-Augmented-Lag (fixed) 26.8 $\pm$ 0.4 -38.6 5312.6 $\pm$ 340.2 758.7 -1219.2 $\pm$ 150.4 -3867.7
MPDO-Robust-Augmented-Lag (linear) 23.5 $\pm$ 0.4 -41.1 6993.3 $\pm$ 325.6 2682.2 -963.4 $\pm$ 122.0 -3066.4
MPDO-Robust-Augmented-Lag (restart) 24.2 $\pm$ 0.5 -41.3 6811.0 $\pm$ 333.5 2581.0 -1043.3 $\pm$ 159.2 -3775.2
MPDO-Robust-Augmented-Lag (geometric) 26.5 $\pm$ 0.5 -40.3 5494.4 $\pm$ 348.4 824.2 -1644.2 $\pm$ 167.2 -4586.5
Table 12: Return and penalised return statistics comparing per-batch schedules for the penalty parameter $\alpha$ under the short runs (200-500 episodes $\times$ maximal number of time steps) and the long runs (2000-5000 episodes $\times$ maximal number of time steps).
Algorithm Return $R_{\text{pen}}^{\pm}$ (signed) $R_{\text{pen}}$ (positive)
Mean $\pm$ SE Min Mean $\pm$ SE Min Mean $\pm$ SE Min
Cartpole (short runs)
MDPO-Robust (fixed) 63.1 $\pm$ 0.1 56.4 -229.7 $\pm$ 179.6 -2610.6 -294.1 $\pm$ 171.4 -2610.8
MDPO-Robust (linear) 50.7 $\pm$ 0.1 44.4 -183.8 $\pm$ 58.8 -831.2 -214.9 $\pm$ 54.7 -833.9
MDPO-Robust (restart) 62.5 $\pm$ 0.1 55.6 -440.3 $\pm$ 173.8 -2450.3 -471.9 $\pm$ 169.0 -2450.5
MDPO-Robust (geometric) 63.1 $\pm$ 0.1 55.6 -286.5 $\pm$ 178.9 -2504.6 -339.3 $\pm$ 171.8 -2504.8
MDPO-Robust-Lag (fixed) 34.4 $\pm$ 0.3 24.9 108.4 $\pm$ 4.4 65.9 34.2 $\pm$ 1.2 24.9
MDPO-Robust-Lag (linear) 35.3 $\pm$ 0.3 25.2 107.1 $\pm$ 5.0 53.0 34.5 $\pm$ 1.2 25.0
MDPO-Robust-Lag (linear) 34.7 $\pm$ 0.3 24.7 108.7 $\pm$ 4.7 62.3 34.4 $\pm$ 1.2 24.7
MDPO-Robust-Lag (restart) 35.4 $\pm$ 0.3 25.1 111.0 $\pm$ 5.1 59.1 35.1 $\pm$ 1.2 25.1
MDPO-Robust-Lag (geometric) 34.6 $\pm$ 0.3 24.3 109.4 $\pm$ 4.7 64.0 34.3 $\pm$ 1.1 24.3
MDPO-Robust-Augmented-Lag (fixed) 34.6 $\pm$ 0.3 24.7 110.8 $\pm$ 4.9 61.6 34.6 $\pm$ 1.2 24.7
MDPO-Robust-Augmented-Lag (linear) 34.5 $\pm$ 0.3 24.4 110.2 $\pm$ 4.7 64.7 34.4 $\pm$ 1.2 24.4
MDPO-Robust-Augmented-Lag (restart) 34.4 $\pm$ 0.3 24.6 108.2 $\pm$ 4.4 67.6 34.0 $\pm$ 1.2 24.6
MDPO-Robust-Augmented-Lag (geometric) 35.0 $\pm$ 0.3 25.0 109.6 $\pm$ 5.1 57.1 34.6 $\pm$ 1.2 25.0
Cartpole (long runs)
MDPO-Robust (fixed) 63.2 $\pm$ 0.1 56.1 -109.7 $\pm$ 176.0 -2647.1 -185.5 $\pm$ 167.1 -2648.3
MDPO-Robust (linear) 18.6 $\pm$ 0.1 15.6 -275.5 $\pm$ 30.8 -630.8 -282.6 $\pm$ 30.8 -640.4
MDPO-Robust (restart) 63.0 $\pm$ 0.1 55.2 -246.4 $\pm$ 174.0 -2510.7 -298.2 $\pm$ 167.1 -2510.9
MDPO-Robust (geometric) 63.1 $\pm$ 0.1 56.2 -151.9 $\pm$ 176.3 -2639.6 -219.7 $\pm$ 168.1 -2640.4
MDPO-Robust-Lag (fixed) 41.6 $\pm$ 0.2 32.1 105.1 $\pm$ 9.5 9.6 38.6 $\pm$ 2.5 2.6
MDPO-Robust-Lag (linear) 22.2 $\pm$ 0.2 16.4 -104.5 $\pm$ 5.2 -165.9 -137.5 $\pm$ 5.1 -199.0
MDPO-Robust-Lag (restart) 41.1 $\pm$ 0.2 30.9 112.4 $\pm$ 8.6 30.1 39.6 $\pm$ 1.7 16.6
MDPO-Robust-Lag (geometric) 41.0 $\pm$ 0.2 30.9 106.3 $\pm$ 8.8 16.2 38.5 $\pm$ 2.2 7.1
MDPO-Robust-Augmented-Lag (fixed) 41.2 $\pm$ 0.2 31.4 106.2 $\pm$ 9.1 13.7 38.6 $\pm$ 2.2 5.0
MDPO-Robust-Augmented-Lag (linear) 25.8 $\pm$ 0.2 19.2 -21.6 $\pm$ 3.3 -46.8 -66.9 $\pm$ 2.8 -102.0
MDPO-Robust-Augmented-Lag (restart) 40.3 $\pm$ 0.2 30.3 111.7 $\pm$ 8.0 33.5 39.0 $\pm$ 1.5 20.0
MDPO-Robust-Augmented-Lag (geometric) 40.9 $\pm$ 0.2 31.1 105.6 $\pm$ 8.8 16.5 38.2 $\pm$ 2.2 6.3
Inventory Management (short runs)
MDPO-Robust (fixed) 118.0 $\pm$ 1.8 72.3 -3306.4 $\pm$ 9.7 -3350.0 -3306.4 $\pm$ 9.7 -3350.0
MDPO-Robust (linear) 120.0 $\pm$ 1.8 73.6 -3425.4 $\pm$ 5.4 -3460.1 -3425.4 $\pm$ 5.4 -3460.1
MDPO-Robust (restart) 116.4 $\pm$ 2.1 55.3 -3240.2 $\pm$ 9.3 -3293.8 -3240.2 $\pm$ 9.2 -3293.8
MDPO-Robust (geometric) 119.0 $\pm$ 1.8 72.3 -3361.8 $\pm$ 7.1 -3382.9 -3361.8 $\pm$ 7.1 -3382.9
MPDO-Robust-Lag (fixed) 60.7 $\pm$ 10.7 -228.5 -930.9 $\pm$ 32.3 -1118.2 -932.2 $\pm$ 30.8 -1118.2
MDPO-Robust-Lag (linear) 56.3 $\pm$ 9.8 -226.8 -939.1 $\pm$ 35.3 -1175.6 -940.5 $\pm$ 33.8 -1175.6
MDPO-Robust-Lag (restart) 68.2 $\pm$ 8.8 -168.2 -964.1 $\pm$ 19.3 -1117.6 -964.5 $\pm$ 18.8 -1117.6
MDPO-Robust-Lag (geometric) 61.9 $\pm$ 8.2 -198.7 -885.0 $\pm$ 24.6 -981.9 -886.9 $\pm$ 21.7 -981.9
MPDO-Robust-Augmented-Lag (fixed) 60.9 $\pm$ 10.6 -221.4 -938.6 $\pm$ 34.4 -1172.2 -940.1 $\pm$ 32.9 -1172.2
MDPO-Robust-Augmented-Lag (linear) 68.9 $\pm$ 8.9 -180.3 -932.6 $\pm$ 28.3 -1118.2 -934.2 $\pm$ 26.7 -1118.2
MDPO-Robust-Augmented-Lag (restart) 66.9 $\pm$ 7.1 -165.4 -837.6 $\pm$ 21.1 -939.1 -838.2 $\pm$ 20.2 -939.1
MDPO-Robust-Augmented-Lag (geometric) 66.8 $\pm$ 7.1 -166.0 -837.3 $\pm$ 21.4 -938.6 -837.9 $\pm$ 20.4 -938.6
Inventory Management (long runs)
MDPO-Robust (fixed) 119.7 $\pm$ 1.8 72.4 -3416.7 $\pm$ 4.1 -3429.7 -3416.7 $\pm$ 4.1 -3429.7
MDPO-Robust (linear) 109.2 $\pm$ 2.9 21.5 -2907.0 $\pm$ 8.5 -2950.6 -2907.1 $\pm$ 8.3 -2950.6
MDPO-Robust (restart) 120.0 $\pm$ 1.8 73.0 -3421.4 $\pm$ 4.7 -3441.4 -3421.4 $\pm$ 4.7 -3441.4
MDPO-Robust (geometric) 114.0 $\pm$ 2.3 50.3 -3121.5 $\pm$ 5.2 -3134.9 -3121.5 $\pm$ 5.2 -3134.9
MPDO-Robust-Lag (fixed) 66.9 $\pm$ 7.1 -166.5 -837.4 $\pm$ 21.4 -939.3 -838.0 $\pm$ 20.4 -939.3
MDPO-Robust-Lag (linear) 60.7 $\pm$ 10.8 -234.4 -930.9 $\pm$ 32.6 -1118.1 -932.4 $\pm$ 30.9 -1118.1
MDPO-Robust-Lag (restart) 66.8 $\pm$ 7.1 -162.6 -837.4 $\pm$ 21.2 -939.0 -837.9 $\pm$ 20.4 -939.0
MDPO-Robust-Lag (geometric) 66.8 $\pm$ 7.1 -163.7 -837.4 $\pm$ 21.3 -939.4 -837.9 $\pm$ 20.5 -939.4
MPDO-Robust-Augmented-Lag (fixed) 66.8 $\pm$ 7.1 -167.7 -837.2 $\pm$ 21.5 -939.1 -837.9 $\pm$ 20.5 -939.1
MDPO-Robust-Augmented-Lag (linear) 66.8 $\pm$ 7.2 -168.8 -837.3 $\pm$ 21.5 -939.1 -837.8 $\pm$ 20.6 -939.1
MDPO-Robust-Augmented-Lag (restart) 66.7 $\pm$ 7.2 -170.1 -837.1 $\pm$ 21.6 -939.5 -837.9 $\pm$ 20.5 -939.5
MDPO-Robust-Augmented-Lag (geometric) 66.7 $\pm$ 7.2 -168.4 -837.2 $\pm$ 21.5 -939.6 -837.9 $\pm$ 20.5 -939.6
3-D Inventory Management (short runs)
MDPO-Robust (fixed) 54.9 $\pm$ 0.4 -9.6 -7956.9 $\pm$ 457.5 -14747.8 -9376.8 $\pm$ 330.2 -14836.3
MDPO-Robust (linear) 55.1 $\pm$ 0.4 -9.5 -8112.0 $\pm$ 455.7 -14726.5 -9507.9 $\pm$ 328.3 -14830.6
MDPO-Robust (restart) 54.5 $\pm$ 0.4 -8.4 -7809.5 $\pm$ 453.2 -14454.0 -9408.3 $\pm$ 318.5 -14561.6
MDPO-Robust (geometric) 55.4 $\pm$ 0.4 -9.3 -8215.7 $\pm$ 454.4 -14973.8 -9588.4 $\pm$ 332.3 -15084.5
MPDO-Robust-Lag (fixed) 41.4 $\pm$ 0.5 -24.6 -1416.9 $\pm$ 446.1 -7726.9 -5088.9 $\pm$ 221.7 -8560.6
MDPO-Robust-Lag (linear) 41.7 $\pm$ 0.5 -26.1 -1573.9 $\pm$ 446.2 -8108.3 -5164.3 $\pm$ 222.7 -8973.5
MDPO-Robust-Lag (restart) 41.0 $\pm$ 0.5 -27.1 -1239.8 $\pm$ 440.8 -7650.5 -5350.6 $\pm$ 217.9 -8969.6
MDPO-Robust-Lag (geometric) 40.2 $\pm$ 0.5 -26.7 -778.5 $\pm$ 441.4 -7152.1 -5115.9 $\pm$ 212.1 -8602.1
MPDO-Robust-Augmented-Lag (fixed) 39.5 $\pm$ 0.5 -29.6 -525.0 $\pm$ 441.2 -6861.9 -4957.0 $\pm$ 211.2 -8289.3
MDPO-Robust-Augmented-Lag (linear) 41.5 $\pm$ 0.5 -25.6 -1474.0 $\pm$ 443.6 -7706.7 -5299.4 $\pm$ 218.6 -8798.9
MDPO-Robust-Augmented-Lag (restart) 40.0 $\pm$ 0.5 -28.7 -754.8 $\pm$ 438.8 -7087.5 -4950.3 $\pm$ 210.5 -8389.0
MDPO-Robust-Augmented-Lag (geometric) 41.6 $\pm$ 0.5 -24.6 -1582.8 $\pm$ 445.8 -8124.9 -5316.9 $\pm$ 218.7 -9080.0
3-D Inventory Management (long runs)
MDPO-Robust (fixed) 87.0 $\pm$ 0.4 24.4 -23312.6 $\pm$ 525.6 -31314.1 -23459.7 $\pm$ 499.3 -31314.1
MDPO-Robust (linear) 100.1 $\pm$ 0.4 41.1 -29655.3 $\pm$ 540.5 -37776.0 -29778.8 $\pm$ 520.6 -37776.0
MDPO-Robust (restart) 93.9 $\pm$ 0.4 32.7 -26709.0 $\pm$ 534.6 -34493.9 -26815.8 $\pm$ 514.0 -34493.9
MDPO-Robust (geometric) 89.8 $\pm$ 0.4 28.8 -24688.1 $\pm$ 520.7 -32646.2 -24771.7 $\pm$ 503.7 -32648.9
MPDO-Robust-Lag (fixed) 32.8 $\pm$ 0.5 -32.4 2617.9 $\pm$ 378.6 -2601.7 -1011.4 $\pm$ 169.9 -3767.4
MDPO-Robust-Lag (linear) 39.3 $\pm$ 0.4 -24.4 -427.1 $\pm$ 372.9 -5798.5 -2588.9 $\pm$ 233.2 -6431.7
MDPO-Robust-Lag (restart) 39.1 $\pm$ 0.4 -27.4 -274.3 $\pm$ 379.8 -5788.0 -2356.2 $\pm$ 228.9 -6272.1
MDPO-Robust-Lag (geometric) 36.7 $\pm$ 0.4 -29.0 602.2 $\pm$ 373.5 -4823.9 -2245.8 $\pm$ 207.3 -5945.4
MPDO-Robust-Augmented-Lag (fixed) 26.8 $\pm$ 0.4 -38.6 5312.6 $\pm$ 340.2 758.7 -1219.2 $\pm$ 150.4 -3867.7
MDPO-Robust-Augmented-Lag (linear) 22.8 $\pm$ 0.5 -43.5 7297.2 $\pm$ 334.4 3001.8 -997.1 $\pm$ 133.8 -3293.2
MDPO-Robust-Augmented-Lag (restart) 22.8 $\pm$ 0.5 -42.8 7290.6 $\pm$ 347.8 2672.1 -679.9 $\pm$ 136.7 -3116.2
MDPO-Robust-Augmented-Lag (geometric) 30.8 $\pm$ 0.4 -33.8 3642.3 $\pm$ 357.3 -1105.7 -1898.7 $\pm$ 200.9 -5429.6